OpenStudy (spitfire77):

If θ is an angle such that cos(θ)= -3/7 and 0 ≤ θ ≤ π, find sin(θ), tan(θ), and sec(θ) if you could demonstrate just one I can do the rest :)

OpenStudy (anonymous):

#IAMTHECALCULASMASTER_

OpenStudy (ddcamp):

If 0 ≤ θ ≤ π, then sin(θ) is positive. Use cos²(θ)+sin²(θ) = 1 to find sin(θ), the rest should follow easily.

OpenStudy (spitfire77):

could you do one of them out for me to go off of?

OpenStudy (ddcamp):

you know cos(θ), so just solve (-3/7)² + sin(θ)² = 1 to find sin(θ).

OpenStudy (anonymous):

\[\theta =\cos ^{-1}\left(-\frac{3}{7}\right)=2.01371 \text{ radians} \]

OpenStudy (anonymous):

sinpower6θ+cospower6θ+3sin^2θcos^2θ=?????????